OpenStudy (anonymous):

how to solve \[2(2xy + 4y -3)dx + (x+2)^2 dy = 0\] via the Integrating Factor Method (Linear Differential Equations)

OpenStudy (saifoo.khan):

@dpaInc

OpenStudy (anonymous):

@Zarkon

OpenStudy (saifoo.khan):

@satellite73

OpenStudy (anonymous):

@turingtest

OpenStudy (anonymous):

2( 2xy + 4y −3 )dx + (x+2)² dy = 0 -> dy/dx = - 4y/(x +2) + 6/ (x+2)² 4y/(x +2) + dy/dx = 6/ (x+2)² Let I = e^ 4∫ dx/(x +2) = e^ [ ln (x+2)^4 ] = ( x+2) ^4 => ( x+2) ^4* y = 6 ∫ ( x+2) ^4 dx/ (x+2)² = 6 ∫ (x+2)² dx

OpenStudy (anonymous):

Can you continue from here?